OpenStudy (anonymous):

What is the series representation of the sequence 4,8,28,80?

OpenStudy (amistre64):

series representation?

OpenStudy (amistre64):

like change from sequences to summation of sequence?

OpenStudy (anonymous):

Yeap!

OpenStudy (amistre64):

i cant see a pattern that would be helpful in determing the sequence

OpenStudy (amistre64):

any ideas?

OpenStudy (anonymous):

i dont have >.< i tried breaking it down by determining the factors. but still, can't see the pattern. :|

OpenStudy (amistre64):

i take it the sequences represents the output of some power series/polynomial ....

OpenStudy (amistre64):

maybe? 4 = c1 8 = c2x 28 = c3(x^2)/2 80 = c4(x^3)/6

OpenStudy (anonymous):

actually the representation i am trying to get is from 4x+8x^2+28x^3+80x^4

OpenStudy (amistre64):

im getting confused in the terminology ...

OpenStudy (anonymous):

summation of sequences* sorry :D

OpenStudy (amistre64):

if i read it right then; you need the "function" that creates these coeffs of this given power series ...

OpenStudy (anonymous):

\[\sum_{n=0}^{\infty} (the sequence of coefficient)\times x ^{n+1}\] this should be the output i just dont know what to put in the coeff

OpenStudy (amistre64):

yep, thats what my thoughts were .... eventually :)

OpenStudy (anonymous):

haha. sorry im really bad at expressing myself :D

OpenStudy (amistre64):

im just wondering if we equate this to the formation of a maclaurin series .... at the moment

OpenStudy (amistre64):

and its just the 4 terms of the sequence right?

OpenStudy (anonymous):

yes though it's up to inf

OpenStudy (anonymous):

my instructor told me to stop the iteration when i see a pattern! hahaha

OpenStudy (amistre64):

4,8,28,80, ... or is the "..." parts = 0?

OpenStudy (anonymous):

yes, that's that!

OpenStudy (amistre64):

whaich part is right lol

OpenStudy (anonymous):

80,.....+n hahaha

OpenStudy (amistre64):

4 8 28 80 4 20 52 16 32 16 cant really say if this makes a good pattern to follow; i dont see enough number to say yes or no to

OpenStudy (anonymous):

80?? 4, 8, 28.. is \[4\sum_{i=1}^{4} i ^{2} - 2\] but 80 is not included

OpenStudy (anonymous):

4 8 28 80 244 728

OpenStudy (amistre64):

.... how many numbers you holding back on ;)

OpenStudy (anonymous):

this is up to infinity!!

OpenStudy (amistre64):

4 8 28 80 244 728 4 20 52 164 484 16 32 112 320 16 80 208 64 128 64 i might see a pattern

OpenStudy (anonymous):

I got something... 1,2,7,20 is each number 3* previous plus minus 1 alternating

OpenStudy (anonymous):

That's after u factor out the 4..

OpenStudy (anonymous):

But that's not quite going to work with 244...

OpenStudy (anonymous):

Oh, yes it does....

OpenStudy (amistre64):

4 8 28 80 244 728 [1532] [5904] any way to know if 4 20 52 164 484 1460 4372 these 2 are next? 16 32 112 320 976 2912 16 80 208 656 1936 64 128 448 1280 64 320 832 256 512 256

OpenStudy (anonymous):

I asked someone but she said that this is not an arithmetic/geometric sequence

OpenStudy (anonymous):

My sequence works....

OpenStudy (anonymous):

can you put it into the algebraic sequences? :D

OpenStudy (anonymous):

Sure, now there is a pattern, it must be possible...

OpenStudy (anonymous):

What's the thing for dealing with an alternating sign, I forgot....

OpenStudy (amistre64):

(-1)^n

OpenStudy (anonymous):

That's it..:-)

OpenStudy (anonymous):

Thanks amistre and estudier. I am going to bed now. Hope I can figure this out by tomorrow morning.